# Nova: Generative Language Models for Assembly Code with Hierarchical Attention and Contrastive Learning

### v4

@@ Abstract
Binary code analysis is the foundation of crucial tasks in the security domain; thus building effective binary analysis techniques is more important than ever. Large language models (LLMs) although have brought impressive improvement to source code tasks, do not directly generalize to assembly code due to the unique challenges of assembly: (1) the low information density of assembly and (2) the diverse optimizations in assembly code. To overcome these challenges, this work proposes a hierarchical attention mechanism that builds attention summaries to capture the semantics more effectively, and designs contrastive learning objectives to train LLMs to learn assembly optimization. Equipped with these techniques, this work develops Nova, a generative LLM for assembly code. Nova outperforms existing techniques on binary code decompilation by up to 146.54%, and outperforms the latest binary code similarity detection techniques by up to 6.17%, showing promising abilities on both assembly generation and understanding tasks.
Nova: Generative Language Models for Assembly Code with
Hierarchical Attention and Contrastive Learning
Nan Jiang1, Chengxiao Wang1,2, Kevin Liu3, Xiangzhe Xu1, Lin Tan1, Xiangyu Zhang1 1Purdue University, 2University of Science and Technology of China, 3Lynbrook High School Correspondence: lintan@purdue.edu
@@ 1 Introduction
Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection Güler et al. (2019); Duan et al. (2020); Chen et al. (2022b), malware detection Spensky et al. (2016); Aonzo et al. (2023); Xu et al. (2014), binary recovery Su et al. (2024); Zhang et al. (2021); Chen et al. (2022c), and legacy software maintenance Carbone et al. (2009); Carlini et al. (2015); Martin et al. (2010). For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand Su et al. (2024); Zhang et al. (2021); Chen et al. (2022c). Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 Sausalito (2020)), effective binary analysis techniques are in high demand.
@@
Large language models pre-trained on source code have brought improvement in various software development domains Chen et al. (2022a); Liu et al. (2023); Chen et al. (2023); Le et al. (2022); Jiang et al. (2023); Xia et al. (2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity Wang et al. (2022); Xu et al. (2023a), malware detection Su et al. (2024), and binary code decompilation Tan et al. (2024); Armengol-Estapé et al. (2024); Hosseini and Dolan-Gavitt (2022). Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs Wang et al. (2022); Su et al. (2024); Xu et al. (2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs. Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs Tan et al. (2024); Armengol-Estapé et al. (2024).
The challenges of leveraging generative LLMs for assembly code are twofold. First, compared to source code, assembly code has a lower information density. A short source-code sequence maps to a much longer assembly-code sequence that is often several times longer. Thus, assembly semantics span across a long sequence of tokens. Figure 1 (a) shows an example of a source code function that compares two integers, while Figure 1 (b) shows its corresponding assembly code optimized with O0 flag. In the O0-optimized assembly code, the five instructions from 10:   mov   -0x8(%rbp),%rax to 1c:   cmp   %eax,%edx perform the checking whether the value of x is smaller than the value of y (correspond to if (*(int*)x < *(int*)y) in the source code). A single assembly instruction alone represents little meaningful semantics in the source code. It is the combinations of many instructions and the dependencies between them represent the semantics. Such combinations of instructions are long, which is hard for LLMs to learn.
@@
  • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly. Nova outperforms state-of-the-art (SOTA) on binary decompilation by up to 146.54% and on binary similarity detection by up to 6.17%.
  • We conduct a comprehensive analysis, illustrating the effectiveness of Nova’s novel designs: (1) Nova’s embeddings of assemblies successfully reflect code functionalities in the latent space, and (2) Nova’s hierarchical attention complements standard attention by learning different attention weight distributions, especially those reflecting long sequence semantics.
@@ 2 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code Rozière et al. (2023); Li et al. (2023); Guo et al. (2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 2.1). Section 2.2 describes Nova’s hierarchical attention. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 2.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 2.4 and 2.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@
2.1 Data Collection
We build our assembly data sets on top of existing source code datasets: The-Stack Li et al. (2023) and the AnghaBench da Silva et al. (2021). We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code. We treat every assembly function as a separate data sample. The breakdown statistics are in Table 1.
@@ 2.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions Vaswani et al. (2023); Radford and Narasimhan (2018); Radford et al. (2019); Brown et al. (2020), our hierarchical self-attention contains three categories at different levels of granularity.
Figure 3: Design of hierarchical attention
@@
(1) Intra-Instruction Attention: Due to the low information density in assembly, intra-instruction attention is designed to capture the summary of every instruction, which is the standard causal attention but limited to tokens of each instruction (the yellow part in Figure 3). Tokens in different instructions have no attention weights. The “[INST]” label at the end of the instruction has attention to all the tokens in the instruction and thus captures the semantics of the entire instruction (e.g., “[INST-1]” captures the semantics of “mov eax, \$1”).
@@
(2) Preceding-Instruction Attention In addition to the local semantics of each instruction, the use of assembly instructions (such as the choice of registers) depends on the context. For example, after the first instruction “mov eax, \$1”, the second instruction should not reuse “eax” to store another value “\$2” immediately. To capture such context, the preceding-instruction attention enables each token in an instruction to have attention to the “[INST]” label of the preceding instruction (the light green part in Figure 3).
@@
(3) Inter-Instruction Attention To understand function semantics (i.e., functionality), which lies in the dependencies across different instructions, the inter-instruction attention is designed to let the “[INST]” label of each instruction have attention to all the labels of previous instructions. For example, “[INST-4]” has attention to “[INST-1]”, “[INST-2]”, and “[INST-3]” (the dark green part in Figure 3). The inter-instruction attention is only enabled for “[INST]” labels, as they represent the semantics of each instruction.
@@
To sum up, the hierarchical self-attention breaks the semantics of assembly code into three parts. The intra-instruction attention captures the instruction summary, and the preceding-instruction attention captures the context with the preceding instruction. The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary. Appendix A.2 shows how hierarchical self-attention works with text and source code.
@@
2.3 Contrastive Learning
The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax. Nova adopts contrastive learning technique Gao et al. (2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics.
The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus Radford and Narasimhan (2018), notated as $L_{lm}$. In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
Functionality CL Functionality CL trains Nova to focus more on the functionalities of assembly code rather than the syntax. Code with the same functionality (assemblies from the same source code), should be encoded closer in the latent space. For instance, in Figure 4 (a), embeddings of source and assembly code of function “cmp” are closer to each other, and the same for function “sort”.
@@ 2.3 Contrastive Learning
$L_{fcl}=-\mathop{log}\sum\limits_{s,t\in S}\sum\limits_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
@@
Optimization CL LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax. Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered.
Higher-level optimization applies a super-set of optimization rules compared to lower-level optimization. Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly. Optimization CL optimizes Nova with the constraint:
$\forall f\in F,\forall s<t_{1}<t_{2}\in S,\;d(e^{s}_{f},e^{t_{1}}_{f})\leq d(e^{s}_{f},e^{t_{2}}_{f})$
Intuitively, this ensures that the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code. For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases. Formally, optimization CL minimizes the following loss:
@@ 2.5 Task 2: Binary Code Similarity Detection
Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets Wang et al. (2022); Su et al. (2024), which is the foundation of various applications such as plagiarism detection Luo et al. (2014); Sæbjørnsen et al. (2009) and vulnerability detection David and Yahav (2014); David et al. (2018, 2017, 2016).
A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$). There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate). Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool. The learning objective is:
@@
, where we follow previous work Su et al. (2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
@@ 3.1 Pre-Training
We use the data collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder Guo et al. (2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers. Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
3.2 Binary Code Decompilation
@@
Training Data We sample 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
@@
Test Data We use HumanEval-Decompile Tan et al. (2024) as the test benchmark, which was not used in training. HumanEval-Decompile is derived from the C language adaptation of the HumanEval Chen et al. (2021) benchmark and provides test cases in evaluating functionality correctness. HumanEval-Decompile contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
Baselines Nova is compared with GPT-3.5, GPT-4, and the previous SOTA LLM4Decompile Tan et al. (2024). LLM4Decompile trains DeepSeek-Coder using the same AnghaBench corpus, and it is the first LLM-based technique that aims to generate executable decompilations.
@@
Evaluation GPT-3.5 and GPT-4 are prompted with three-shot examples, while LLM4Decompile and Nova samples 20 decompilations per assembly, using the temperature of 0.2 and top_p of 0.95 Chen et al. (2021). The generated decompilations are executed against the test cases and both Pass@1 and Pass@10 Chen et al. (2021) are reported.
3.3 Binary Code Similarity Detection

### v5

@@ Abstract
Binary code analysis is the foundation of crucial tasks in the security domain; thus building effective binary analysis techniques is more important than ever. Large language models (LLMs) although have brought impressive improvement to source code tasks, do not directly generalize to assembly code due to the unique challenges of assembly: (1) the low information density of assembly and (2) the diverse optimizations in assembly code. To overcome these challenges, this work proposes a hierarchical attention mechanism that builds attention summaries to capture the semantics more effectively, and designs contrastive learning objectives to train LLMs to learn assembly optimization. Equipped with these techniques, this work develops Nova, a generative LLM for assembly code. Nova outperforms existing techniques on binary code decompilation by up to 14.84 – 21.58% (absolute percentage point improvement) higher Pass@1 and Pass@10, and outperforms the latest binary code similarity detection techniques by up to 6.17% Recall@1, showing promising abilities on both assembly generation and understanding tasks.
@@ 1 Introduction
Binary code plays an irreplaceable role in the security domain, being the foundation of crucial tasks including vulnerability detection (Güler et al., 2019; Duan et al., 2020; Chen et al., 2022b), malware detection (Spensky et al., 2016; Aonzo et al., 2023; Xu et al., 2014), binary recovery (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c), and legacy software maintenance (Carbone et al., 2009; Carlini et al., 2015; Martin et al., 2010). For example, when performing tasks such as identifying attacks and malware, security analysts often only have access to assembly, i.e., the human-readable representation of binary code, which is extremely difficult to understand (Su et al., 2024; Zhang et al., 2021; Chen et al., 2022c). Thus, combined with the increasing sophistication of cybercrime that poses significant threats worldwide (e.g., cybercrime is predicted to cost the world \$10.5 trillion annually by 2025 (Sausalito, 2020)), effective binary analysis techniques are in high demand.
@@
Large language models pre-trained on source code have brought improvement in various software development domains (Chen et al., 2022a; Liu et al., 2023; Chen et al., 2023; Le et al., 2022; Jiang et al., 2023; Xia et al., 2023). However, these LLMs are not designed for or trained with assembly corpus, not achieving their full potential on binary code analysis tasks such as binary code similarity (Wang et al., 2022; Xu et al., 2023a), malware detection (Su et al., 2024), and binary code decompilation (Tan et al., 2024; Armengol-Estapé et al., 2024; Hosseini & Dolan-Gavitt, 2022).
Existing work applying LLMs on assembly code mainly piggybacks on encoder-style LLMs (Wang et al., 2022; Su et al., 2024; Xu et al., 2023a), unable to benefit from the more extensive pre-training, updated architectures, scaling of state-of-the-art generative LLMs. Other work using generative LLMs for decompilation shows a low unit test passing rate of the decompiled programs (Tan et al., 2024; Armengol-Estapé et al., 2024).
The challenges of leveraging generative LLMs for assembly code are twofold. First, compared to source code, assembly code has a lower information density. A short source-code sequence maps to an assembly-code sequence that is often several times longer. Thus, assembly semantics span across a long sequence of tokens. Figure 1 (a) shows an example of a source code function that compares two integers, while Figure 1 (b) shows its corresponding assembly code optimized with O0 flag. In the O0-optimized assembly code, the five instructions from 10:   mov   -0x8(%rbp),%rax to 1c:   cmp   %eax,%edx perform the checking whether the value of x is smaller than the value of y (correspond to if (*(int*)x < *(int*)y) in the source code). A single assembly instruction alone represents little meaningful semantics in the source code. It is the combinations of many instructions and the dependencies between them represent the semantics. Such combinations of instructions are long, which is hard for LLMs to learn.
@@
  • We develop Nova, a generative foundation LLM with hierarchical attention and contrastive learning for assembly. Nova outperforms state-of-the-art on binary decompilation by up to 14.84 – 21.58% higher Pass@k and on binary similarity detection by up to 6.17% Recall@1.
@@ 2 Approach
Figure 2 presents the overall approach of Nova. We build Nova on top of foundation models for source code (Guo et al., 2024) to utilize their source code and natural language generation ability. We first collect large assembly corpora (Section 2.1). Section 2.2 describes Nova’s hierarchical attention design. With the collected assembly corpora, we then pretrain Nova with language modeling and contrastive learning objectives (Section 2.3). Then, we fine-tune Nova on two important downstream tasks, binary code decompilation, and binary code similarity detection (Sections 2.4 and 2.5), to prove Nova’s effectiveness and benefits to the binary research domain.
@@
2.1 Data Collection
We build our assembly data sets on top of existing source code datasets: The-Stack (Li et al., 2023) and the AnghaBench (da Silva et al., 2021). We compile the source code into executables with different optimization levels (i.e., O0, O1, O2 and O3), strip the executables to remove debug information, and disassemble them into assembly code. We treat every assembly function as a separate data sample. The breakdown statistics are in Table 2.
@@ 2.2 Hierarchical Self-Attention
Nova uses hierarchical self-attention that is specially designed to learn the low-information-density semantics in the long sequence of assembly code. Specifically, Nova learns the assembly code in an hierarchical way by providing a modified attention mask. Different from standard token-level attentions (Vaswani et al., 2023; Radford & Narasimhan, 2018; Radford et al., 2019; Brown et al., 2020), our hierarchical self-attention contains three categories at different levels of granularity.
Figure 3: Design of Nova’s hierarchical attention for assembly code
@@
(1) Intra-Instruction Attention: Due to the low information density in assembly, intra-instruction attention is designed to capture the summary of every instruction, which is the standard causal attention but limited to tokens of each instruction (the yellow part in Figures 3) (a) and (b). Tokens in different instructions have no attention weights. The “[INST]” label at the end of the instruction has attention to all the tokens in the instruction and thus captures the semantics of the entire instruction (e.g., “[INST-1]” captures the semantics of “mov eax, \$1”).
@@
(2) Preceding-Instruction Attention: In addition to the local semantics of each instruction, the use of assembly instructions (such as the choice of registers) depends on the context. For example, after the first instruction “mov eax, \$1”, the second instruction should not reuse “eax” to store another value “\$2” immediately. To capture such context, the preceding-instruction attention enables each token in an instruction to have attention to the “[INST]” label of the preceding instruction (the light green part in Figures 3 (a) and (b)).
@@
(3) Inter-Instruction Attention: To understand function semantics (i.e., functionality), which lies in the dependencies across different instructions, the inter-instruction attention is designed to let the “[INST]” label of each instruction have attention to all the labels of previous instructions. For example, “[INST-4]” has attention to “[INST-1]”, “[INST-2]”, and “[INST-3]” (the dark green part in Figures 3 (a) and (b)). The inter-instruction attention is only enabled for “[INST]” labels, as they represent the semantics of each instruction.
@@
To sum up, the hierarchical self-attention breaks the semantics of assembly code into three parts. The intra-instruction attention captures the instruction summary, and the preceding-instruction attention captures the context with the preceding instruction. The inter-instruction attention learns the long dependencies across instructions on top of the “[INST]” labels that contain the instruction summary. Figure 3 (c) also shows how Nova’s hierarchical attention for assembly code is compatible with the standard self-attention for text and source code. As existing LLMs have shown good performance on text and source code using the standard self-attention, we keep the standard causal attention mask within and between any chunks of text or source code in the input (the light grey part shown in Figure 3 (c)). The attention from text or source code to assembly code (and vice versa) is only allowed for the “[INST]” tokens as they are designed to contain the assembly instruction summaries.
2.3 Contrastive Learning
The syntax gap between assembly code and source code, and syntax diversity between differently-optimized assembly code make LLMs struggle to distinguish the semantics behind the syntax. Nova adopts contrastive learning technique (Gao et al., 2021) during pre-training to train LLMs to encode assembly code meaningfully w.r.t semantics. The standard pre-training objective is language modeling by minimizing the negative likelihood of code in the pre-training corpus (Radford & Narasimhan, 2018), notated as $L_{lm}$. In addition, Nova is pre-trained with two new objectives, $L_{fcl}$ for functionality contrastive learning and $L_{ocl}$ for optimization contrastive learning.
@@ 2.3 Contrastive Learning
Functionality CL: Functionality CL trains Nova to focus more on the functionalities of assembly code rather than the syntax. Code with the same functionality (assemblies from the same source code), should be encoded closer in the latent space. For instance, in Figure 4 (a), embeddings of source and assembly code of function “cmp” are closer to each other, and the same for function “sort”.
@@
$L_{fcl}=-\mathop{log}\sum_{s,t\in S}\sum_{f_{i}\in F}\left(1-\frac{\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{i}})\right)}{\sum_{f_{j}\in F}\mathop{exp}\left(d(e^{s}_{f_{i}},e^{t}_{f_{j}})\right)}\right)$
@@
Optimization CL: LLMs can be confused if being asked to directly connect a source code function to its O3-optimized assembly, due to their dramatically different syntax. Such a huge gap can be filled by learning how the source code is transformed to O0, O1, O2 and eventually to O3 assembly, as the optimization levels are ordered.
Higher-level optimization applies a super-set of optimization rules compared to lower-level optimization. Nova learns such order with the optimization CL objective, encoding differently-optimized assembly code orderly. Optimization CL optimizes Nova with the constraint: the more optimizations applied, the larger the difference between embeddings of optimized and unoptimized code. For instance, Figure 4 (c) and (d) illustrate that for the same function “cmp”, the distance between source code and assembly increases when the optimization level increases. Formally, optimization CL minimizes the following loss:
@@ 2.5 Task 2: Binary Code Similarity Detection
Binary code similarity detection (BCSD) aims to measure the similarity between two binary code snippets (Wang et al., 2022; Su et al., 2024), which is the foundation of various applications such as plagiarism detection (Luo et al., 2014; Sæbjørnsen et al., 2009) and vulnerability detection (David & Yahav, 2014; David et al., 2018; 2017; 2016).
A widely used setting is taking a query assembly of the function $f^{q}$ that is compiled with one optimization level (denoted by $s$), and a pool of candidate assembly of $K$ functions (notated by $f^{p}_{i},$ $1\leq i\leq K$) compiled with a different optimization level (denoted by $t\neq s$). There exists a unique candidate assembly coming from the same source code as the query ($\exists!1\leq i\leq K,$ $f^{p}_{i}=f^{q}$, called the positive candidate). Nova is fine-tuned to encode these binaries, so that the positive candidate has the highest similarity with the query assembly among the pool. The learning objective is as follows:
@@
We follow previous work (Su et al., 2024) to let $s$ be O0-assembly and $t$ be O3-assembly, which is the hardest setting.
@@ 3 Experimental Setup
This section describes the setup of pre-training and fine-tuning of Nova, as well as the existing baselines we compare Nova with, and the evaluation metrics we used in the two downstream tasks. Appendix A.2 contains additional details such as training hyper-parameters.
@@ 3.1 Pre-Training
We use the data collected from AnghaBench and The-Stack for pre-training. We pre-train Nova starting from DeepSeek-Coder (Guo et al., 2024), and the hierarchical attention is applied on half of the attention heads to balance between its effectiveness and the existing knowledge in the standard attention layers (Justified in Appendix A.3). Nova is pre-trained with language modeling for one epoch, followed by contrastive learning objectives for another epoch.
3.2 Fine-Tuning for Binary Code Decompilation
@@
Training Data: We sample (due to computation resource limitation) 2.16M assembly-to-source-code pairs (0.338B tokens) from the pre-training corpus to build the BCD fine-tuning data.
@@
Test Data: We use HumanEval-Decompile (Tan et al., 2024) as the test benchmark, which was not used in training. HumanEval-Decompile is derived from the C language adaptation of the HumanEval (Chen et al., 2021) benchmark and contains 164 C functions, each compiled with O0O3 optimization flags and disassembled into X86-64 assembly.
Baselines: Nova is compared with existing SOTA LLM4Decompile (Tan et al., 2024), Meta LLMCompiler (Cummins et al., 2024), and other general code LLMs (CodeLlama (Rozière et al., 2023), StarCoder2 (Lozhkov et al., 2024), DeepSeekCoder (Guo et al., 2024), Qwen2.5-Coder (Hui et al., 2024), Granite Mishra et al. (2024), LongCoder (Guo et al., 2023), GPT-3.5-Turbo, and GPT-4o). LLM4Decompile trains DeepSeekCoder using the same AnghaBench corpus for binary decompilation. Meta LLMCompiler trains CodeLlama models using LLVM IRs, X86, and ARM assembly code to optimize and translate binary code.
@@
Evaluation: GPT-3.5-Turbo and GPT-4o are prompted with three-shot examples, while all the other LLMs and Nova sample 20 decompilations per assembly code, using the temperature of 0.2 and top_p of 0.95 (Chen et al., 2021). The generated decompilations are executed against the test cases. Finally, Pass@1 and Pass@10 (Chen et al., 2021) are reported.
3.3 Fine-Tuning for Binary Code Similarity Detection